the formula 1/R = 1/R1 + 1/R2, concerning resistance in an electrical circuit is used in physics. if R = 3 ohms and R1 is 8 ohms more than R2, the value of R1 be?

this also not make any sense to me
Answer #1
R2 = R1 - 8
R1 R2 = R R2 + R R1 = R(R1+R2)

R1(R1-8) = 3(2 R1 - 8)

R1^2 - 8 R1 = 6 R1 - 24

R1^2 -14 R1 + 24 = 0

(R1-12)(R1-2) = 0
If R1 were 2, R2 would be negative, no good
so
R1 = 12 (and R2 is 4)
answered by: chalbi
Answer #2
that make sense now :) thanks very much damon :)
answered by: Tyshera